OpenStudy (anonymous):

. Joey has a beach ball with a radius of 20 cm. What volume of air will it hold when fully inflated? A) 83.8 cm3 B) 251.4 cm3 C) 1676.2 cm3 D) 33,510.3 cm3

OpenStudy (anonymous):

do you know the formula of the volume of a sphere ? ?

OpenStudy (anonymous):

=4/3*3.14*20^3 V=4.1867*8,000 V=33,493 cm^3 ans. is d

OpenStudy (anonymous):

no i dont i think its (4/3) pi r 3

OpenStudy (anonymous):

i think b

OpenStudy (anonymous):

which one ? b or d ? im confused now

OpenStudy (anonymous):

robert2 1=4/3*3.14*20^3 V=4.1867*8,000 V=33,493 is d